### v5

# Partition Regularity of Nonlinear Rado Systems: Necessary and Sufficient Conditions

Abstract
A matrix $A$ is called kernel partition regular if, for every finite coloring of the natural numbers $\mathbb{N}$, there exists a monochromatic solution to the equation $A\vec{X}=0$. In 1933, Rado characterized such matrices by showing that a matrix is kernel partition regular if and only if it satisfies the so-called column condition. In this article, we investigate polynomial extensions of Rado’s theorem. We exhibit several nonlinear systems of equations that are kernel partition regular and demonstrate that satisfying the column condition continues to ensure kernel partition regularity even when the system is augmented with a polynomial term. Then we prove a necessary condition for the partition regularity of Nonlinear Rado systems.
Mathematics subject classification 2020: 05D10, 05C55, 22A15, 54D35.
Keywords: Nonlinear partition regularity, Rado’s column condition, Kernel partition regularity, Central Sets, Ramsey theory.
1   Introduction
Van der Waerden  achieved one of the earliest breakthroughs in combinatorial number theory citevdw, who proved that every finite coloring of the natural numbers contains arbitrarily long arithmetic progressions in at least one color class. Rado significantly extended this foundational result [18], who provided a complete characterization of homogeneous linear systems that admit monochromatic solutions under any finite partition of $\mathbb{N}$. In this article, we use the term partition regular to indicate that a certain structure can be found monochromatically under any finite coloring of $\mathbb{N}$. For instance, a collection $\mathcal{F}$ of subsets of $\mathbb{N}$ is said to be partition regular if, for every finite coloring of $\mathbb{N}$, there exists a set $F\in\mathcal{F}$ that lies entirely within a single color class. Similarly, an equation of the form $F(x_{1},\ldots,x_{n})=0$ over $\mathbb{N}$ is said to be partition regular if, for every finite coloring of $\mathbb{N}$, the equation admits a monochromatic solution.
To state Rado’s classification theorem, we first recall the definition of the column condition.
Definition 1.1 (Column Condition).
Let $m,n\in\mathbb{N}$, and let $A$ be an $m\times n$ matrix with integer entries. Denote the columns of $A$ by $(C_{1},\ldots,C_{n})$. The matrix $A$ is said to satisfy the column condition if there exists a partition of the index set $\{1,2,\ldots,n\}$ into subsets $I_{1}\cup\cdots\cup I_{v}$ such that:
  • $\sum_{j\in I_{1}}C_{j}=\vec{0}$, and
  • for each $t>1$, the sum $\sum_{j\in I_{t}}C_{j}$ can be expressed as a rational linear combination of the columns $\{C_{j}:j\in I_{1}\cup\cdots\cup I_{t-1}\}$.
A matrix $A$ is called kernel partition regular if, for every finite coloring of $\mathbb{N}$, the equation $A\vec{X}=0$ admits a monochromatic solution. Rado’s theorem asserts that a matrix $A$ is kernel partition regular if and only if it satisfies the column condition. For some recent development on kernel partition regularity we refer to the articles [1, 14].
To obtain an arithmetic progression of length $m$, one considers the matrix
$A=\begin{pmatrix}1&1&-1&0&\cdots&0\\ 1&2&0&-1&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 1&m&0&0&\cdots&-1\\ \end{pmatrix},$
which corresponds to the statement of van der Waerden’s theorem. Another classical example is the matrix (1  11)111(1\;\;1\;\;-1)( 1 1 - 1 ), which satisfies the column condition and encodes one of the earliest results in Ramsey theory: Schur’s theorem[19]. This theorem asserts that the pattern $\{x,y,x+y\}$ is partition regular. By applying the map $n\mapsto 2^{n}$, one immediately obtains the multiplicative analogue of Schur’s theorem, which states that the pattern $\{x,y,x\cdot y\}$ is partition regular. In [5], Csikvári, Gyarmati, and Sárközy conjectured that the equation $a+b=c\cdot d$ is partition regular. This conjecture was independently resolved by Bergelson and Hindman in [3, 12], using tools from the algebraic structure of the Stone–Čech compactification of discrete semigroups. In this article, we study polynomial generalizations of two classical results: Rado’s characterization of kernel partition regularity and the theorem asserting the partition regularity of the equation $a+b=c\cdot d$.
The study of polynomial generalizations of classical results in Ramsey theory is inspired by [4, 11], where the authors established the polynomial van der Waerden theorem. Let $\mathbb{P}$ denote the set of polynomials with rational coefficients and zero constant term. The polynomial van der Waerden theorem states that for any finite subset $F\subset\mathbb{P}$, and for any finite coloring of $\mathbb{N}$, there exist $a,d\in\mathbb{N}$ such that the set $\{a+P(d):P\in F\}$ is monochromatic.
The first goal of this article is to demonstrate the existence of certain nonlinear systems of equations that are partition regular. To achieve this, we extend the sum-equals-product theorem to a polynomial setting. In addition, we establish an infinitary version of the sum-equals-product theorem. Our approach involves a novel technique using a special class of ultrafilters that are simultaneously large with respect to both addition and multiplication [^2] A detailed discussion of these ultrafilters is provided later.. This method enables us to fuse additive polynomial structures with multiplicatively large sets, ultimately yielding the desired results.
For example, one of our results shows that for any polynomial $P\in\mathbb{P}$, the following system of equations is partition regular:
$x\cdot y$ $=z_{1}+P(z)$
$x\cdot y^{2}$ $=z_{2}+P(z)$
Another example of our approach reveals that the following equation is partition regular.
$x_{1}+x_{2}$ $=y_{1}\cdot z$
$x_{1}+2x_{2}$ $=y_{2}\cdot z.$
To extend Rado’s result in the nonlinear settings, we introduce a class of nonlinear systems.
Definition 1.2 (Nonlinear Rado System).
Let $A$ be an $m\times n$ matrix (with $n\geq 2$) with integer coefficients that satisfies the column condition, and let $P_{1},\ldots,P_{m}$ be elements of $\mathbb{P}$, the set of rational-coefficient polynomials with zero constant term. Consider the system of equations:
$a_{1,1}x_{1}+\cdots+a_{1,n-1}x_{n-1}+a_{1,n}\cdot y_{1}+P_{1}(z)$ $=0$
$\vdots$
$a_{m,1}x_{1}+\cdots+a_{m,n-1}x_{n-1}+a_{m,n}\cdot y_{m}+P_{m}(z)$ $=0$
This system is called a Nonlinear Rado System.
Note that, in contrast to the linear case where one typically sets $y_{1}=\cdots=y_{m}$, here we allow the variables $y_{1},\ldots,y_{m}$ to be distinct. This distinction is necessary due to the nonlinear terms $P_{i}(z)$. However, if all the polynomials $P_{i}$ are identically zero, then the system reduces to the linear case, and we may take $y_{1}=\cdots=y_{m}$. We can summarize certain nonlinear Rado systems using a matrix framework. Given any matrix $A=(a_{i,j})$ of size $m\times n$, we define the expanded matrix of $A$ as follows:
$E(A)=\begin{pmatrix}a_{1,1}&\cdots&a_{1,n-1}&a_{1,n}&0&\cdots&0\\ a_{2,1}&\cdots&a_{2,n-1}&0&a_{2,n}&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ a_{m,1}&\cdots&a_{m,n-1}&0&0&\cdots&a_{m,n}\\ \end{pmatrix}.$
If $P_{1},P_{2},\ldots,P_{m}$ are elements of $\mathbb{P}$, then we define the column matrix $P$ as
$P=\begin{pmatrix}P_{1}\\ P_{2}\\ \vdots\\ P_{m}\end{pmatrix}.$
Then, the system of equations described in Definition 1.2 can be expressed compactly as
$E(A)\vec{X}+P(z)=0.$
Our second main result establishes that every nonlinear Rado system is partition regular. In other words, if the matrix $A$ satisfies the column condition, then the equation
$E(A)\vec{X}+P(z)=0$
is partition regular.
Towards a Necessary Condition
Our third major result concerns identifying a necessary condition under which a system of polynomial equations remains partition regular. Let $A=(a_{i,j})$ be an $m\times l$ matrix, and let $P=\{P_{1},\ldots,P_{m}\}\subset\mathbb{P}$ be a fixed finite set of polynomials. Let $n$ denote the maximum degree among the polynomials in $P$. For any $z\in\mathbb{N}$, the vector $P(z)\in\mathbb{Z}^{m}$ can be expressed as:
$P(z)=\begin{pmatrix}a_{1,l+1}\\ a_{2,l+1}\\ \vdots\\ a_{m,l+1}\end{pmatrix}z+\begin{pmatrix}a_{1,l+2}\\ a_{2,l+2}\\ \vdots\\ a_{m,l+2}\end{pmatrix}z^{2}+\cdots+\begin{pmatrix}a_{1,l+n}\\ a_{2,l+n}\\ \vdots\\ a_{m,l+n}\end{pmatrix}z^{n}.$
Here, for each $1\leq i\leq m$, the polynomial $P_{i}(z)$ takes the form
$P_{i}(z)=a_{i,l+1}z+a_{i,l+2}z^{2}+\cdots+a_{i,l+n}z^{n},$
where some of the coefficients may be zero.
Define
$C=\begin{pmatrix}a_{1,l+1}&a_{1,l+2}&\cdots&a_{1,l+n}\\ a_{2,l+1}&a_{2,l+2}&\cdots&a_{2,l+n}\\ \vdots&\vdots&\cdots&\vdots\\ a_{m,l+1}&a_{m,l+2}&\cdots&a_{m,l+n}\end{pmatrix}$
to be the coefficient matrix associated with the polynomial vector $P$, and let
$A_{\text{aug}}(P)=\begin{pmatrix}A&C\end{pmatrix}$
be the augmented matrix formed by adjoining $A$ with the coefficients of $P$.
Define the submatrix
$A_{\text{aug}}^{\text{lin}}(P)=\begin{pmatrix}a_{1,1}&a_{1,2}&\cdots&a_{1,l+1}\\ a_{2,1}&a_{2,2}&\cdots&a_{2,l+1}\\ \vdots&\vdots&\cdots&\vdots\\ a_{m,1}&a_{m,2}&\cdots&a_{m,l+1}\end{pmatrix}$
to be the part of $A_{\text{aug}}(P)$ corresponding to the linear terms of the polynomials in $P$.
The following theorem provides a necessary condition for the partition regularity of the nonlinear Rado system.
Before stating our main result, define $c_{t_{0}}=\vec{0}$, and for each $11\leq j\leq n-$ 1, let $c_{t_{j}}=c_{l+j+1}$.
Theorem 1.3.
Let $P=\{P_{1},\ldots,P_{m}\}\subset\mathbb{P}$ be a finite set of polynomials, and let $A=(a_{ij})$ be an $m\times l$ matrix with rational entries. Suppose that for every finite coloring of $\mathbb{N}$, there exists a monochromatic set $\vec{X}=\{x_{1},\ldots,x_{l}\}\cup\{z\}\subset\mathbb{N}$ such that the equation
$A\vec{X}+P(z)=\vec{0}$
is satisfied. Then there exists $s\in\mathbb{N}$ and a partition $[1,l+1]=\cup_{j=1}^{s}I_{j}$ such that the following conditions hold:
  1. 1.
    We have that $\sum_{i\in I_{1}}c_{i}=0$; that is, there exists a subset $I_{1}\subseteq\{1,2,\ldots,l+1\}$ such that the sum of the columns of $A_{\text{aug}}^{\text{lin}}(P)$ indexed by $I_{1}$ is equal to the zero vector; and
  2. 2.
    For each $1\leq j\leq s$, there exists $10\leq k\leq n-$ 1 such that
    $\sum_{i\in I_{j}}c_{i}\in\operatorname{span}_{\mathbb{Q}}\left(\{c_{p}:p\in\cup_{k=1}^{j-1}I_{k}\}\cup\{c_{t_{i}}:0\leq i\leq k\}\right).$
  3. 3.
    There exists $0\leq q\leq s$ such that
    $c_{l+2}\in\operatorname{span}_{\mathbb{Q}}\{c_{i}:i\in\cup_{k=1}^{q}I_{k}\}.$
  4. 4.
    For each $3\leq u\leq n$, there exists $0\leq v\leq s$ such that
    $c_{l+u}\in\operatorname{span}_{\mathbb{Q}}\left(\{c_{i}:i\in\cup_{k=1}^{v}I_{k}\}\cup\{c_{l+2},\ldots,c_{l+u-1}\}\right).$
Note that if each $P_{i}$ is chosen to be the zero polynomial, then the theorem immediately reduces to the classical column condition of Rado. Furthermore, if each $P_{i}$ is a linear polynomial with zero constant term, i.e., $P_{i}(z)=c_{i}z$ for some constants $c_{i}$, then the system reduces to the linear form $A_{\text{aug}}^{\text{lin}}(P)\vec{X}=\vec{0}$, and the result asserts that $A_{\text{aug}}^{\text{lin}}(P)$ satisfies the column condition.
By multiplying the entire system $A\vec{X}+P(z)=\vec{0}$ by a sufficiently large integer, one can ensure that all entries of the matrix $A$ and all coefficients of the polynomials $P_{i}(z)$ become integers. Therefore, it suffices to prove the above theorem under the assumption that both the matrix $A$ and the coefficients of $P(z)$ have integer entries.
The study of necessary conditions for nonlinear Rado systems traces back to the work of Bergelson and Hindman in [3, 12], where they proved that the nonlinear equation $a+b-cd=0$ is partition regular. This naturally led to the question of identifying necessary conditions that any system of nonlinear partition regular equations must satisfy.
Note that in [6], the authors studied necessary conditions for certain nonlinear equations, while in [2], the authors introduced the notion of the maximal Rado condition and established a necessary condition for a single nonlinear equation. Both these works rely on tools from nonstandard analysis. In contrast, our result provides a necessary condition for an entire system of nonlinear equations using purely combinatorial arguments.
Our proofs rely on the machinery of ultrafilters. Before proceeding to the main results, we briefly recall some fundamental concepts related to ultrafilters and their connection to Ramsey theory.
2   Preliminaries of Ultrafilters
Ultrafilters are set-theoretic objects that play a central role in modern Ramsey theory. In this section, we give a brief overview of the basic notions. For a comprehensive treatment, we refer the reader to the book by Hindman and Strauss [16].
A filter $\mathcal{F}$ on a nonempty set $X$ is a collection of subsets of $X$ satisfying the following properties:
  1. 1.
    $\emptyset\notin\mathcal{F}$ and $X\in\mathcal{F}$,
  2. 2.
    If $A\in\mathcal{F}$ and $A\subseteq B\subseteq X$, then $B\in\mathcal{F}$,
  3. 3.
    If $A,B\in\mathcal{F}$, then $A\cap B\in\mathcal{F}$.
Using Zorn’s Lemma, one can show that every filter is contained in a maximal filter, called an ultrafilter. Any ultrafilter $p$ on $X$ satisfies the following partition property:
  • For every finite partition $X=A_{1}\cup\cdots\cup A_{r}$, there exists $i\in\{1,2,\ldots,r\}$ such that $A_{i}\in p$.
Let $S$ be a discrete semigroup. The elements of $\beta S$, the Stone-Čech compactification of $S$, are regarded as ultrafilters on $S$. For any subset $A\subseteq S$, we define the set $\overline{A}=\{p\in\beta S:A\in p\}$. The collection $\{\overline{A}:A\subseteq S\}$ forms a basis for the closed sets in $\beta S$. The semigroup operation $\cdot$ on $S$ extends to the Stone-Čech compactification $\beta S$, making $(\beta S,\cdot)$ a compact right topological semigroup. This means that for each $p\in\beta S$, the function $\rho_{p}(q):\beta S\to\beta S$, defined by $\rho_{p}(q)=q\cdot p$, is continuous, and $S$ is contained within the topological center of $\beta S$. Specifically, for any $x\in S$, the function $\lambda_{x}:\beta S\to\beta S$, defined by $\lambda_{x}(q)=x\cdot q$, is also continuous.
This structure is connected to a famous result by Ellis, which states that if $S$ is a compact right topological semigroup, then the set of idempotents $E(S)\neq\emptyset$. In semigroup theory, a nonempty subset $I$ of a semigroup $T$ is called a left ideal of $S$ if $TI\subset I$, a right ideal if $IT\subset I$, and a two-sided ideal (or simply an ideal) if it is both a left and right ideal. A minimal left ideal is a left ideal that does not contain any proper left ideal, and similarly, we can define a minimal right ideal and the smallest ideal.
For any compact Hausdorff right topological semigroup $T$, the smallest two-sided ideal, denoted $K(T)$, is given by:
$K(T)=\bigcup\{L:L\text{ is a minimal left ideal of }T\}=\bigcup\{R:R\text{ is a minimal right ideal of }T\}.$
Given a minimal left ideal $L$ and a minimal right ideal $R$, their intersection $L\cap R$ is a group, and in particular, it contains an idempotent. If $p$ and $q$ are idempotents in $T$, we write $p\leq q$ if and only if $pq=qp=p$. An idempotent is minimal with respect to this relation if and only if it is a member of the smallest ideal $K(T)$.
For $p,q\in\beta S$ and $A\subseteq S$, it holds that $A\in p\cdot q$ if and only if the set $\{x\in S:x^{-1}A\in q\}\in p$, where $x^{-1}A=\{y\in S:x\cdot y\in A\}$.
$\bullet$ A set $A\subseteq S$ is said to be a piecewise syndetic if there exists $p\in K(\beta S,\cdot)$ such that $A\in p.$
Ellis’ Theorem asserts that every compact right topological semigroup contains an idempotent. If $T$ is a compact right topological semigroup, we denote by $E(T)$ the set of all idempotents of $T$, i.e.,
$E(T):=\{e\in T:e\cdot e=e\}.$
The existence of idempotents in $\beta\mathbb{N}$ is intimately related to Hindman’s Theorem [10]. For any nonempty set $X$, let $\mathcal{P}_{f}(X)$ denote the collection of all nonempty finite subsets of $X$. For any commutative semigroup $(S,+)$, a set $A\subseteq S$ is said to be an IP set if there exists a sequence $\langle x_{n}\rangle_{n}$ in $S$ such that
$A=\operatorname{FS}(\langle x_{n}\rangle_{n})=\left\{\sum_{t\in H}x_{t}:H\in\mathcal{P}_{f}(\mathbb{N})\right\}.$
Similarly, for any $r\in\mathbb{N}$, a set $A\subseteq S$ is called an IPr set if there exists a sequence $\langle x_{n}\rangle_{n=1}^{r}$ such that
$A=\operatorname{FS}(\langle x_{n}\rangle_{n=1}^{r})=\left\{\sum_{t\in H}x_{t}:\emptyset\neq H\subseteq\{1,2,\ldots,r\}\right\}.$
From the ultrafilter proof of Hindman’s Theorem [^3] This is an unpublished result due to Galvin and Glazer, it follows that if $A\in p$ for some idempotent ultrafilter $p$, then $A\subseteq S$ is an IP set. Moreover, this condition is also sufficient; that is, $A$ is an IP set if and only if $A\in p$ for some idempotent $p\in\beta S$.
Definition 2.1.
A set $A\subseteq S$ is central if there exists $p\in E\left(K(\beta S,\cdot)\right)$ such that $A\in p.$
A set $A\subseteq\mathbb{N}$ is called additively large (respectively, multiplicatively large) if it is large in the semigroup $(\mathbb{N},+)$ (respectively, $(\mathbb{N},\cdot)$).
From [4, 11], it follows that every additively piecewise syndetic set witnesses the conclusion of the polynomial van der Waerden theorem.
It is a standard exercise to verify that the set $\overline{E(\beta\mathbb{N},+)}$, the closure of the idempotents in $(\beta\mathbb{N},+)$, forms a left ideal in the semigroup $(\beta\mathbb{N},\cdot)$. Consequently, the intersection
$\overline{E(\beta\mathbb{N},+)}\cap E(\beta\mathbb{N},\cdot)$
is nonempty.
Similarly, the closure of the minimal ideal $\overline{E(K(\beta\mathbb{N},+))}$ is a left ideal in $(\beta\mathbb{N},\cdot)$, which implies that
$\overline{E(K(\beta\mathbb{N},+))}\cap E(K(\beta\mathbb{N},\cdot))$
is also nonempty.
Any ultrafilter $p$ in this intersection is called a centrally rich ultrafilter. If $p$ is centrally rich and $A\in p$, then $A$ is both additively and multiplicatively central.
Before we proceed to our main proof, let us recall one final interesting fact about ultrafilters, which we will use in our proof. A set $A\subseteq\mathbb{N}$ is called an $IP_{0}$ set if, for every $r\in\mathbb{N}$, the set $A$ contains an $IP_{r}$ set. Clearly, every member of an additive idempotent ultrafilter is an $IP_{0}$ set.
We denote by $\overline{\mathrm{IP}_{0}}$ the set of all ultrafilters in $\beta\mathbb{N}$ whose every member is an $IP_{0}$ set. It is a routine exercise to verify that $\overline{\mathrm{IP}_{0}}$ is a two-sided ideal in the semigroup $(\beta\mathbb{N},\cdot)$. In particular, this implies that
$\overline{K(\beta\mathbb{N},\cdot)}\subseteq\overline{\mathrm{IP}_{0}}.$
3   Proof of Our Main Results
Extending the Sum-Equals-Product Theorem
In this section, we focus on extending the classical Sum-Equals-Product phenomenon in Ramsey theory. We explore both infinitary and polynomial extensions of this idea, showing that even complex combinations of sums and products can exhibit partition regularity.
Infinitary Extension of the Sum-Equals-Product Theorem
We begin by establishing an infinitary version of the Sum-Equals-Product phenomenon. Our goal is to construct sequences whose finite sums and finite products, as well as certain structured combinations of the two, lie entirely within a single cell of any finite coloring of the natural numbers. This result significantly strengthens earlier partition regularity results and lays the foundation for subsequent polynomial generalizations.
Theorem 3.1.
Let $r\in\mathbb{N}$, and suppose that $\mathbb{N}=\bigcup_{i=1}^{r}A_{i}$ is a finite partition. Then there exist two sequences $\langle a_{n}\rangle_{n}$ and $\langle b_{n}\rangle_{n}$ in $\mathbb{N}$, and an index $i\in\{1,2,\ldots,r\}$, such that:
  1. 1.
    $FS\left(\langle a_{n}\rangle_{n}\right)\cup FP\left(\langle b_{n}\rangle_{n}\right)\subseteq A_{i}$, and
  2. 2.
    For every $N\in\mathbb{N}$, we have
    $\bigcup_{m=1}^{N}FS\left(\langle a_{n}\rangle_{n=1}^{m}\right)\cdot FP\left(\langle b_{n}\rangle_{n=m}^{N}\right)\subseteq A_{i}.$
  • Since $(a_{1}+a_{2})\cdot b_{2}=a_{1}\cdot b_{2}+a_{2}\cdot b_{2}$, the theorem above implies that the equation $a+b=c\cdot d$ is partition regular. That is, no matter how the natural numbers are finitely colored, there exists a monochromatic solution to the equation $a+b=c\cdot d$.
  • Also, observe that
    $(a_{1}+a_{2}+\cdots+a_{n})\cdot b_{n+1}b_{n+2}\cdots b_{n+(m-1)}=\sum_{i=1}^{n}a_{i}\cdot b_{n+1}b_{n+2}\cdots b_{n+(m-1)}.$
    This shows that, for any $m,n\in\mathbb{N}$, the equation
    $x_{1}+x_{2}+\cdots+x_{n}=y_{1}\cdot y_{2}\cdots y_{m}$
    is partition regular, as proved in [12].
Proof of Theorem 3.1.
Choose any
$p\in\overline{E\left(\beta\mathbb{N},+\right)}\cap E\left(\beta\mathbb{N},\cdot\right),$
and let $A\in p=p\cdot p$. Define:
$A^{\star}=\{x\in A:x^{-1}A\in p\}\in p.$
Then $A^{\star}$ is an additive IP-set. Hence, there exists $q\in E(\beta\mathbb{N},+)$ such that $A^{\star}\in q$.
Now define:
$A^{\star\star}=\{x\in A^{\star}:-x+A^{\star}\in q\}\in q.$
Then, by [16, Lemma 4.14], for all $x\in A^{\star\star}$, we have:
$-x+A^{\star\star}\subseteq A^{\star\star}.$
Choose $a_{1}\in A^{\star\star}$. Since $A^{\star\star}\subseteq A^{\star}$, we also have $a_{1}\in A^{\star}$. Then the intersection $A^{\star}\cap a_{1}^{-1}A^{\star}\in p$, because both sets belong to $p$, and $p$ is closed under finite intersections.
Now choose:
$b_{1}\in A^{\star}\cap a_{1}^{-1}A^{\star}.$
Then:
$a_{1}\in A^{\star},\quad b_{1}\in A^{\star},\quad\text{and}\quad a_{1}\cdot b_{1}\in A^{\star}.$
So,
$\{a_{1},b_{1},a_{1}\cdot b_{1}\}\subseteq A^{\star}\subseteq A.$
Now, since $a_{1}\in A^{\star\star}$, and by Lemma 4.14 from [16], we have:
$A^{\star\star}\cap(-a_{1}+A^{\star\star})\in q.$
So we choose:
$a_{2}\in A^{\star\star}\cap(-a_{1}+A^{\star\star}).$
This implies:
$\{a_{1},a_{2},a_{1}+a_{2}\}\subseteq A^{\star\star}\subseteq A^{\star}.$
Now define:
$b_{2}\in A^{\star}\cap\bigcap_{x\in\{a_{1},a_{2},a_{1}+a_{2}\}}x^{-1}A^{\star}\cap\bigcap_{y\in\{a_{1},b_{1},a_{1}\cdot b_{1}\}}y^{-1}A^{\star}.$
Then:
$\{b_{2},a_{1}\cdot b_{2},a_{2}\cdot b_{2},(a_{1}+a_{2})\cdot b_{2},a_{1}\cdot b_{1}\cdot b_{2}\}\subseteq A^{\star}\subseteq A.$
Hence, we have the following inclusions:
  1. 1.
    $\{a_{1},a_{2},a_{1}+a_{2}\}\subset A^{\star\star}\subseteq A^{\star},$
  2. 2.
    $\{b_{2}\}\cup b_{2}\cdot\{a_{1},a_{2},a_{1}+a_{2}\}\subseteq A^{\star},$
  3. 3.
    $\{a_{1},b_{1},a_{1}\cdot b_{1}\}\subset A^{\star},$
  4. 4.
    $b_{2}\cdot\{a_{1},b_{1},a_{1}\cdot b_{1}\}\subset A^{\star}.$
Inductively, we assume that for some $N\in\mathbb{N}$, we have two sequences $\langle a_{n}\rangle_{n=1}^{N}$ and $\langle b_{n}\rangle_{n=1}^{N}$ such that:
  1. 1.
    $FS\left(\langle a_{n}\rangle_{n=1}^{N}\right)\subset A^{\star\star},$
  2. 2.
    $FP\left(\langle b_{n}\rangle_{n=1}^{N}\right)\subset A^{\star},$
  3. 3.
    For every $M\leq N$, we have
    $\bigcup_{m=1}^{M}FS\left(\langle a_{n}\rangle_{n=1}^{m}\right)\cdot FP\left(\langle b_{n}\rangle_{n=m}^{M}\right)\subset A^{\star}.$
To complete the induction, choose
$a_{N+1}\in A^{\star\star}\cap\bigcap_{x\in FS\left(\langle a_{n}\rangle_{n=1}^{N}\right)}(-x+A^{\star\star})\in q.$
Hence,
$FS\left(\langle a_{n}\rangle_{n=1}^{N+1}\right)\subset A^{\star\star}.$
Define
$B=\bigcup_{M=1}^{N}\bigcup_{m=1}^{M}FS\left(\langle a_{n}\rangle_{n=1}^{m}\right)\cdot FP\left(\langle b_{n}\rangle_{n=m}^{M}\right)\subset A^{\star}.$
Now choose
$b_{N+1}\in A^{\star}\cap\bigcap_{x\in FS\left(\langle a_{n}\rangle_{n=1}^{N}\right)}x^{-1}A^{\star}\cap\bigcap_{y\in B}y^{-1}A^{\star}\cap\bigcap_{z\in FP\left(\langle b_{n}\rangle_{n=1}^{N}\right)}z^{-1}A^{\star}\in p.$
Now clearly we have:
  1. 1.
    $FS\left(\langle a_{n}\rangle_{n=1}^{N+1}\right)\subset A^{\star\star}$,
  2. 2.
    $FP\left(\langle b_{n}\rangle_{n=1}^{N+1}\right)\subset A^{\star}$, and
  3. 3.
    $FS\left(\langle a_{n}\rangle_{n=1}^{N+1}\right)\cdot b_{N+1}\cup B\cdot b_{N+1}\subseteq A^{\star}$.
Now, from the definition of $B$, it is clear that this verifies the third inductive hypothesis for $N+1$ as well.
This completes the proof. ∎
However the technique of the proof of the above theorem can be generalized further to prove a more general result. Before we describe that let us introduce two notions of regular structure.
Definition 3.2.
  1. 1.
    (Additive regular): A family of finite subsets $\mathcal{F}$ of $\mathbb{N}$ is called additive regular if every additive central set contains an element $F\in\mathcal{F}.$
  2. 2.
    (Multiplicative regular): A family of finite subsets $\mathcal{G}$ of $\mathbb{N}$ is called multiplicative regular if every multiplicative central set contains an element $G\in\mathcal{G}.$
If we closely look at the proof of Theorem 3.1, we see that one can replace each $a_{n}$ by an element of a given additive regular family, and each $b_{n}$ by an element of a given multiplicative regular family.
For any two finite sets $F,G$, define
$F\cdot G=\{f\cdot g:f\in F,\,g\in G\}.$
Given a sequence of finite sets $\langle F_{n}\rangle_{n=1}^{N}$, define:
  • $FS\left(\langle F_{n}\rangle_{n=1}^{N}\right)=\bigcup_{f_{i}\in F_{i}\text{ for each }i\in\{1,\ldots,N\}}FS\left((f_{i})_{i=1}^{N}\right),$
  • $FP\left(\langle F_{n}\rangle_{n=1}^{N}\right)=\bigcup_{f_{i}\in F_{i}\text{ for each }i\in\{1,\ldots,N\}}FP\left((f_{i})_{i=1}^{N}\right).$
The proof of the following theorem is similar to the proof of Theorem 3.1, and hence we omit it.
Theorem 3.3.
Let $\langle\mathcal{F}_{n}\rangle_{n}$ and $\langle\mathcal{G}_{n}\rangle_{n}$ be sequences of additive and multiplicative regular families, respectively.
Then for any $r\in\mathbb{N}$ and any finite partition $\mathbb{N}=\bigcup_{i=1}^{r}A_{i}$, there exist sequences $\langle F_{n}\rangle_{n}$ and $\langle G_{n}\rangle_{n}$ such that:
  1. 1.
    For every $n\in\mathbb{N}$, we have $F_{n}\in\mathcal{F}_{n}$ and $G_{n}\in\mathcal{G}_{n}$;
  2. 2.
    $FS\left(\langle F_{n}\rangle_{n}\right)\cup FP\left(\langle G_{n}\rangle_{n}\right)\subseteq A_{i}$ for some $i\in\{1,2,\ldots,r\}$;
  3. 3.
    For every $N\in\mathbb{N}$,
    $\bigcup_{m=1}^{N}FS\left(\langle F_{n}\rangle_{n=1}^{m}\right)\cdot FP\left(\langle G_{n}\rangle_{n=m}^{N}\right)\subseteq A_{i}.$
Corollary 3.4.
For any $l,m,n\in\mathbb{N}$ with $n>2$, $m>1$, the system of nonlinear equations
$x_{1}+x_{2}+x_{3}+\cdots+x_{n}$ $=z_{1}\cdot y_{1}y_{2}\cdots y_{m},$
$x_{1}+2x_{2}+x_{3}+\cdots+x_{n}$ $=z_{2}\cdot y_{1}y_{2}\cdots y_{m},$
$\vdots$
$x_{1}+lx_{2}+x_{3}+\cdots+x_{n}$ $=z_{l}\cdot y_{1}y_{2}\cdots y_{m},$
is partition regular.
Proof.
Define the families:
  • $\mathcal{F}_{1}=\{\{a,a+d,a+2d,\dots,a+ld\}:a,d\in\mathbb{N}\}$,
  • $\mathcal{F}_{2}=\left\{\left\{b_{1}+\cdots+b_{n-2}\right\}:b_{i}\in\mathbb{N}\text{ for all }i\right\}$,
  • $\mathcal{G}_{2}=\left\{\left\{c_{1}\cdots c_{m-1}\right\}:c_{i}\in\mathbb{N}\right\}$.
By Theorem 3.3, there exist sets $F_{1}\in\mathcal{F}_{1}$, $F_{2}\in\mathcal{F}_{2}$, and $G_{2}\in\mathcal{G}_{2}$ such that $(F_{1}+F_{2})\cdot G_{2}$ is monochromatic.
Then, for each $i\in\{1,\dots,l\}$, the quantity
$(a+id+b_{1}+\cdots+b_{n-2})\cdot c_{1}\cdots c_{m-1}$
belongs to $(F_{1}+F_{2})\cdot G_{2}$, and can be rewritten as
$x_{1}+ix_{2}+x_{3}+\cdots+x_{n}=y_{1}\cdots y_{m}.$
This gives a monochromatic solution for each row in the system. Hence, the system is partition regular. ∎
In the above proof, we may choose $\mathcal{G}_{2}$ to be the set of geometric progressions of length $m$ to obtain the following corollary. The proof is similar and thus omitted.
Corollary 3.5.
For any $l,m,n\in\mathbb{N}$ with $n>2$ and $m>1$, the system of nonlinear equations
$x_{1}+x_{2}+x_{3}+\cdots+x_{n}$ $=y_{1}\cdot z^{m},$
$x_{1}+2x_{2}+x_{3}+\cdots+x_{n}$ $=y_{2}\cdot z^{m},$
$\vdots$
$x_{1}+lx_{2}+x_{3}+\cdots+x_{n}$ $=y_{l}\cdot z^{m},$
is partition regular.
Polynomial Extension of Sum-Equals-Product Theorems
We now extend the sum-equals-product phenomenon to a broader class of nonlinear systems involving polynomial perturbations. Specifically, we show that even when additive expressions are equated to products modified by polynomials, the system remains partition regular.
Theorem 3.6 (Polynomial Sum-Equals-Product Theorem).
Let $m,n,r\in\mathbb{N}$, and let $\{P_{i}:i=1,2,\ldots,r\}\subset\mathbb{P}$ be a finite collection of polynomials with zero constant term. Then the following system of nonlinear equations is partition regular:
$x_{1}+x_{2}+\cdots+x_{n}$ $=y_{1}y_{2}\cdots y_{m}\cdot z_{1}+P_{1}(y_{m+1}),$
$x_{1}+x_{2}+\cdots+x_{n}$ $=y_{1}y_{2}\cdots y_{m}\cdot z_{2}+P_{2}(y_{m+1}),$
$\vdots$
$x_{1}+x_{2}+\cdots+x_{n}$ $=y_{1}y_{2}\cdots y_{m}\cdot z_{r}+P_{r}(y_{m+1}).$
Proof.
We begin by assuming that $m\geq 2$.
Let $p$ be a centrally rich ultrafilter and let $A\in p$. Since $p=p\cdot p$, it follows that the multiplicative star set
$A^{\star}:=\{x\in A:x^{-1}A\in p\}$
also belongs to $p$.
A well-known fact from ultrafilter theory asserts that
$\overline{K(\beta\mathbb{N},\cdot)}\subseteq\overline{\mathrm{IP}_{0}},$
so every member of $p$ contains finite products of some IP-sequence. In particular, since $A^{\star}$ is multiplicatively piecewise syndetic, there exists a finite sequence $\langle a_{i}\rangle_{i=1}^{n}\subset\mathbb{N}$ such that
$FS\left(\langle a_{i}\rangle_{i=1}^{n}\right)\subset A^{\star}.$
Define the set
$B:=A^{\star}\cap\bigcap_{x\in FS\left(\langle a_{i}\rangle_{i=1}^{n}\right)}x^{-1}A^{\star}.$
Then $B\in p$, and so $B^{\star}\in p$ as well.
Now choose a finite sequence $\langle b_{j}\rangle_{j=1}^{m-1}$ such that
$FP\left(\langle b_{j}\rangle_{j=1}^{m-1}\right)\subset B^{\star}.$
Define
$C:=B^{\star}\cap\bigcap_{y\in FP\left(\langle b_{j}\rangle_{j=1}^{m-1}\right)}y^{-1}B^{\star}.$
Since $C\in p$ and $p$ is centrally rich, it follows that $C$ is an additively central set.
By the polynomial van der Waerden theorem, we can choose $a,d\in\mathbb{N}$ such that
$\left\{a,\,d,\,a-\frac{1}{(a_{1}+\cdots+a_{n})\cdot b_{1}\cdots b_{m-1}}\cdot P_{i}(d):i=1,2,\ldots,r\right\}\subset C.$
Thus,
$\left\{(a_{1}+\cdots+a_{n})\cdot b_{1}\cdots b_{m-1}\cdot\left(a-\frac{1}{(a_{1}+\cdots+a_{n})\cdot b_{1}\cdots b_{m-1}}\cdot P_{i}(d)\right):i=1,2,\ldots,r\right\}\subset A^{\star}.$
Now define:
  1. 1.
    For each $i\in\{1,\ldots,r\}$,
    $z_{i}:=a-\frac{1}{(a_{1}+\cdots+a_{n})\cdot b_{1}\cdots b_{m-1}}\cdot P_{i}(d),$
  2. 2.
    For each $j\in\{1,\ldots,n\}$,
    $x_{j}:=a_{j}\cdot b_{1}\cdots b_{m-1}\cdot a,$
  3. 3.
    For each $k\in\{1,\ldots,m\}$,
    • If $k\neq m$, then $y_{k}:=b_{k}$,
    • If $k=m$, then $y_{m}:=a_{1}+\cdots+a_{n}$,
  4. 4.
    Finally, let $y_{m+1}:=d$.
Now, for each $i\in\{1,\ldots,r\}$, we compute:
$(a_{1}+\cdots+a_{n})\cdot b_{1}\cdots b_{m-1}\cdot\left(a-\frac{1}{(a_{1}+\cdots+a_{n})\cdot b_{1}\cdots b_{m-1}}\cdot P_{i}(d)\right)$
$=(a_{1}+\cdots+a_{n})\cdot b_{1}\cdots b_{m-1}\cdot a-P_{i}(d)$
$=x_{1}+\cdots+x_{n}-P_{i}(y_{m+1}).$ (1)
But again,
$(a_{1}+\cdots+a_{n})\cdot b_{1}\cdots b_{m-1}\cdot\left(a-\frac{1}{(a_{1}+\cdots+a_{n})\cdot b_{1}\cdots b_{m-1}}\cdot P_{i}(d)\right)$
$=(a_{1}+\cdots+a_{n})\cdot b_{1}\cdots b_{m-1}\cdot z_{i}$
$=y_{1}\cdots y_{m}\cdot z_{i}.$ (2)
Equating expressions (1) and (2), we conclude that:
$x_{1}+\cdots+x_{n}=y_{1}\cdots y_{m}\cdot z_{i}+P_{i}(y_{m+1}),$
as required.
Now for each $i\in\{1,\ldots,r\}$, from equations (1) and (2), we have
$x_{1}+\cdots+x_{n}=y_{1}\cdots y_{m}\cdot z_{i}+P_{i}(y_{m+1}).$
For the case $m=1$, we interpret the product $b_{1}\cdots b_{m-1}$ as the empty product, i.e., equal to 1. Then the entire construction and argument go through verbatim.
This completes the proof. ∎
Note 3.7.
Note that in Theorem 3.6, we can rename $y_{1}\cdots y_{m}z_{i}$ as $z_{i}$ to conclude that the following system of equations is partition regular:
$x_{1}+\cdots+x_{n}$ $=z_{1}+P_{1}(z),$
$\vdots$
$x_{1}+\cdots+x_{n}$ $=z_{r}+P_{r}(z).$
In addition, for every subset $F\subseteq\{1,2,\ldots,n\}$, the sum $\sum_{t\in F}x_{t}$ also lies in the same cell of the partition.
Theorem 3.8.
The following result is a polynomial extension of the Sum-Equals-Product theorem in a different direction. We adapt the technique of the proof of Theorem 3.6 to prove it.
Let $n\in\mathbb{N}$, and let $\{P_{i}:i=1,2,\ldots,n\}\subset\mathbb{P}$ be a collection of polynomials. Then the system of equations
$x\cdot y$ $=z_{1}+P_{1}(z),$
$x\cdot y^{2}$ $=z_{2}+P_{2}(z),$
$\vdots$
$x\cdot y^{n}$ $=z_{n}+P_{n}(z)$
is partition regular. In addition, the set $\{xy^{i}:1\leq i\leq n\}$ also lies entirely within a single cell of the partition.
Proof.
Let $p$ be a centrally rich ultrafilter and $A\in p$. Then $A^{\star}\in p$. As $A^{\star}$ is multiplicatively central, from [tran, Corollary 2.7], there exist $a,d\in\mathbb{N}$ such that
$F=\{a,d,ad,ad^{2},\ldots,ad^{n}\}\subset A^{\star}.$
Define
$B=A^{\star}\cap\bigcap_{s\in F}s^{-1}A^{\star}\in p.$
Now $B$ is additively central (since $p$ is centrally rich), so there exist $b,c\in\mathbb{N}$ such that
$\left\{b,c,b-\frac{1}{s}P_{i}(c):s\in F,\ i=1,2,\ldots,n\right\}\subset B.$
Now define:
  1. 1.
    $y=d$;
  2. 2.
    $z=c$;
  3. 3.
    $x=ab$.
Then for every $i=1,2,\ldots,n$, we define:
$z_{i}$ $=ad^{i}\left(b-\frac{1}{ad^{i}}P_{i}(c)\right)$
$=xy^{i}-P_{i}(z).$
Thus, each equation
$xy^{i}=z_{i}+P_{i}(z)$
is satisfied with all variables lying in the same color class. This completes the proof. ∎
The following corollary of the above theorem shows that rational functions involving polynomials are also partition regular.
Corollary 3.9.
Let $n\in\mathbb{N}$, and $P,Q\in\mathbb{P}$. Then the equation
$\frac{x+P(d)}{y+Q(d)}=z^{n}$
is partition regular.
In fact, for any $P,Q,R,S\in\mathbb{P}$, the equation
$\frac{w+P(d)}{x+Q(d)}=\left(\frac{y+R(d)}{z+S(d)}\right)^{n}$
is partition regular.
Nonlinear Rado Systems are Partition Regular
Now we prove that nonlinear Rado systems, as defined earlier, are indeed partition regular.
Theorem 3.10.
Nonlinear Rado systems are partition regular.
Proof.
Let $A=(a_{i,j})$ be an $m\times n$ matrix satisfying the conditions of Definition 1.2. Let $p$ be a centrally rich ultrafilter. Since $p\in K(\beta\mathbb{N},\cdot)$, every set in $p$ contains a solution $\vec{X}\in\mathbb{N}^{n}$ to the homogeneous linear system $A\vec{X}=\vec{0}$.
Fix $B\in p$, and define
$B^{\star}:=\{n\in\mathbb{N}:n^{-1}B\in p\}\in p.$
Then $B^{\star}\in p$, and there exists $\vec{X}=\begin{bmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{n}\end{bmatrix}\in(B^{\star})^{n}$ such that $A\vec{X}=\vec{0}$.
Now define
$C:=B^{\star}\cap\bigcap_{i=1}^{n}x_{i}^{-1}B^{\star}\in p.$
Let $P_{1},\ldots,P_{m}\in\mathbb{P}$ be given polynomials. Since $C$ is additively central, there exist $a,d\in\mathbb{N}$ such that
$\left\{a,d,\ a-\frac{1}{a_{i,n}x_{n}}P_{i}(d):1\leq i\leq m\right\}\subseteq C.$
Now for each $i=1,2,\ldots,m$, compute:
$a_{i,1}ax_{1}+\cdots+a_{i,n-1}ax_{n-1}+a_{i,n}x_{n}\left(a-\frac{1}{a_{i,n}x_{n}}P_{i}(d)\right)$
$=a_{i,1}ax_{1}+\cdots+a_{i,n-1}ax_{n-1}+a_{i,n}ax_{n}-P_{i}(d)$
$=a\left(a_{i,1}x_{1}+\cdots+a_{i,n}x_{n}\right)-P_{i}(d)$
$=0-P_{i}(d)=-P_{i}(d).$
Thus, each row of the nonlinear Rado system is satisfied, and all relevant elements lie in the same color class determined by $p$. Hence, the system is partition regular. ∎
Example 3.11.
As a direct consequence of the previous theorem, the following nonlinear system of equations is partition regular:
$x_{1}+2x_{2}-3y_{1}+(z^{2}+z)=0,$
$0\displaystyle 2x_{1}-x_{2}-y_{2}+z^{3}=$ 0 .
4   Proof of the Necessary Condition
In this section, we prove Theorem 1.3. Our proof adapts the classical argument of Rado, as found in [8, 17]. The main idea involves the use of the $p$-adic representation of natural numbers. Before presenting the full proof, we require the following preparatory lemma.
Lemma 4.1.
([8, Lemma 6][17, Lemma 4 ]) Let $j,k\in\mathbb{N}$, and let $c_{1},\dots,c_{j}\in\mathbb{Z}^{k}$ be vectors such that $c_{1}$ is not a rational linear combination of the (possibly empty) set $\{c_{2},\dots,c_{j}\}$. Then there exists a finite set of primes $F$ such that for any prime $p\notin F$ and any non-negative integer $n$, the vector $p^{n}c_{1}$ is not a linear combination of $c_{2},\dots,c_{j}\mod p^{n+1}$.
Now we start our proof.
Proof of Theorem 1.3.
Let $P=\{P_{1},\ldots,P_{m}\}\subset\mathbb{P}$ be a finite set of polynomials, and let $A=(a_{ij})\in\mathbb{Z}^{m\times l}$ be an integer matrix. Suppose that for every finite coloring of $\mathbb{N}$, there exists a monochromatic set $\vec{X}=\{x_{1},\ldots,x_{l}\}\cup\{z\}$ such that the system
$A\vec{X}+P(z)=\vec{0}$
holds.
Let $c_{1},\dots,c_{l+n}$ denote the columns of the augmented matrix $A_{\mathrm{aug}}(P)$. Consider any two disjoint subsets $I,J\subset\{c_{1},\dots,c_{l+n}\}$ with $I$ nonempty. Then, either the sum $\sum_{c\in I}c$ lies in the $\mathbb{Q}$-linear span of $J$, or Lemma 4.1 applies and yields a finite set of primes $F_{I,J}$ satisfying its conclusion. Define $F$ to be the (finite) union of all such sets $F_{I,J}$, taken over all choices of disjoint subsets $I,J$ with $I\neq\emptyset$. Let $p$ be any prime not belonging to $F$.
Let $\pi:\mathbb{Z}\to\mathbb{Z}/(p\mathbb{Z})$ denote the natural projection. Define a coloring $\chi_{p}:\mathbb{N}\to\{1,\dots,p-1\}$ recursively as follows:
$\chi_{p}(x)=\begin{cases}\pi(x)&\text{if }\pi(x)\neq 0,\\ \pi(x/p)&\text{if }\pi(x)=0.\end{cases}$
Let $(x_{1},\dots,x_{l+1})\in\mathbb{N}^{l+1}$ be a monochromatic vector such that, defining $\vec{X}=(x_{1},\dots,x_{l})$ and $z=x_{l+1}$, we obtain a solution to the equation $A\vec{X}+P(z)=\vec{0}$. Then there exists some $a\in\{1,\dots,p-1\}$ such that for each $i$, the congruence $x_{i}\equiv ap^{n_{i}}\mod p^{n_{i}+1}$ holds for some $n_{i}\in\{0,1,\dots\}$.
Let $s+1$ be the number of distinct values in the set $\{n_{i}:i=1,\dots,l+1\}$. After reordering the columns of $A$, if needed, we can choose indices $10=d_{0}<d_{1}<d_{2}<\cdots<d_{s}<d_{s+1}=l+$ 1 such that
$n_{i}=\begin{cases}b_{0}&\text{if }d_{0}<i\leq d_{1},\\ b_{1}&\text{if }d_{1}<i\leq d_{2},\\ \vdots&\vdots\\ b_{s}&\text{if }d_{s}<i\leq d_{s+1},\end{cases}$
for some strictly increasing sequence $b_{0}<b_{1}<\cdots<b_{s}$.
For each $11\leq i\leq s+$ 1, define the index set $I_{i}=[d_{i-1}+1,d_{i}]$. Then
$0=\sum_{i\in I_{1}}x_{i}c_{i}=ap^{b_{0}}\sum_{i\in I_{1}}c_{i}\pmod{p^{b_{0}+1}}.$
Since higher-degree terms (such as $x_{l+1}^{2},x_{l+1}^{3},\dots$) do not contribute modulo $p^{b_{0}+1}$, the computation of the sum over $I_{1}$ is straightforward. However, for subsequent sums, complications may arise if $x_{l+1}$ belongs to $I_{1}$. To avoid these issues, we introduce the following modification.
For the general case, first define $c_{t_{0}}=\vec{0}$. Now for every $j>1$, there exists $l+2=t_{1}\leq t_{k}\leq l+n$ such that
00 0 $=\sum_{i=1}^{l+1}x_{i}c_{i}+\sum_{i=2}^{n}c_{l+i}x_{l+1}^{i}$
$\equiv\sum_{i=1}^{d_{j}}x_{i}c_{i}+ap^{b_{j}}\sum_{i=d_{j}+1}^{d_{j+1}}c_{i}+\sum_{i=0}^{k}c_{t_{i}}x_{l+1}^{i+1}\pmod{p^{b_{j}+1}}.$
Hence,
$(-a)p^{b_{j}}\sum_{i\in I_{j+1}}c_{i}\equiv\sum_{i\in I_{1}\cup\cdots\cup I_{j}}x_{i}c_{i}+\sum_{i=0}^{k}c_{t_{i}}x_{l+1}^{i+1}\pmod{p^{b_{j}+1}}.$
Let $\tilde{a}\in\mathbb{Z}$ be an integer such that $a\tilde{a}\equiv 1\pmod{p^{b_{j}+1}}$. (Such an integer exists because $a$ is coprime to $p$, and hence also to $p^{b_{j}+1}$.) Then we obtain the following congruence:
$p^{b_{j}}\sum_{i\in I_{j+1}}c_{i}\equiv-\tilde{a}\sum_{i\in I_{1}\cup\cdots\cup I_{j}}x_{i}c_{i}-\tilde{a}\sum_{i=0}^{k}c_{t_{i}}x_{l+1}^{i+1}\pmod{p^{b_{j}+1}}.$ (3)
Again, suppose $x_{l+1}\in I_{j+1}$ for some $0\leq j\leq s$. Then $x_{l+1}=ap^{b_{j}}$. Let $2\leq u\leq n$. Then there exists $0\leq k\leq s$ such that
00 0 $=\sum_{i=1}^{l+1}x_{i}c_{i}+\left(c_{l+2}x_{l+1}^{2}+\cdots+c_{l+n}x_{l+1}^{n}\right)$
$=\sum_{i=1}^{d_{k+1}}x_{i}c_{i}+c_{l+2}x_{l+1}^{2}+\cdots+c_{l+u}x_{l+1}^{u}\pmod{p^{ub_{j}+1}}$
$\equiv\sum_{i=1}^{d_{k+1}}x_{i}c_{i}+a^{2}c_{l+2}p^{2b_{j}}+\cdots+a^{u}c_{l+u}p^{ub_{j}}\pmod{p^{ub_{j}+1}}.$
Again, since $(a,p)=1$, we let $\tilde{a}$ denote the inverse of $a^{u}$ modulo $p^{ub_{j}+1}$. Then we obtain the congruence
$p^{ub_{j}}c_{l+u}\equiv-\tilde{a}\left(\sum_{i=1}^{d_{k+1}}x_{i}c_{i}+a^{2}c_{l+2}p^{2b_{j}}+\cdots+a^{u-1}c_{l+u-1}p^{(u-1)b_{j}}\right)\pmod{p^{ub_{j}+1}}.$ (4)
Since $p\notin F$, it follows from the definition of $F$ that the previous congruences—equations (3) and (4)—imply the following:
$c_{d_{j}+1}+c_{d_{j}+2}+\cdots+c_{d_{j+1}}\in\operatorname{span}_{\mathbb{Q}}\left(\{c_{i}:i\leq d_{j}\}\cup\{c_{t_{i}}:0\leq i\leq k\}\right),$
and for each $2\leq u\leq n$,
$c_{l+u}\in\operatorname{span}_{\mathbb{Q}}\left(\{c_{i}:i\leq d_{k}\}\cup\{c_{l+1},c_{l+2},\dots,c_{l+u-1}\}\right).$
5   Further Possibilities
Our current treatment is restricted to systems with finitely many equations and variables. However, infinite linear systems have been studied extensively (see [14]). A natural extension would be to ask:
Is it possible to identify infinite nonlinear Rado systems that remain partition regular?
In [13, 15], the authors studied monochromatic solutions to systems of linear equations of the form $A\vec{X}=\vec{b}$. They showed that such a system is partition regular—that is, it admits a monochromatic solution under every finite coloring of $\mathbb{N}$—if and only if it has a constant solution; that is, there exists $d\in\mathbb{Q}$ such that
$A\begin{pmatrix}d\\ d\\ \vdots\\ d\end{pmatrix}=\vec{b}.$
A natural question arises:
Can one formulate a nonlinear analogue of this result?
Acknowledgement
I sincerely thank Dibyendu De for his valuable suggestions on the previous draft of this paper. I also gratefully acknowledge the support of the NBHM Postdoctoral Fellowship (Reference No: 0204/27/(27)/2023/R & D-II/11927).